• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jack (Techs)

    Log entry time 16:17:51 on October 26, 2009

    Entry number 296716

    keyword=Status of spectrometers

    Problems with the spectrometers have a few suspected causes. Four of the liquid level meters have failed. It is unlikely the failures are linked to background radiation. More likely it would seem to be fallout from the power loss on Tuesday. Two meters on each arm have failed. We have some basis for suspecting the internal power supplies of the meters. Ongoing radiation damage to the massflow meters on the leadflows is causing them to drift. This makes adjustments difficult to keep track of. We have set the leadflows on the Left Arm to what seems to be a usable setting. Serious changes in the cryo flows could vary the leadflows and cause a trip on the lead voltage, but if things are stable this should not be a problem. One of the valve controllers on the Right Dipole had an issue. A CAMAC module was replaced and control was restored. The RQ1 trip at 15:40 yesterday seems to have been due to a failed Newport process controller. The dump of RQ1 then proceeded to bring everything else down. Since the experiment only needs one working spectrometer to finish some calibrations, we elected to bring the Left Arm up as soon as possible to ensure the calibrations could be done. We have turned off the magnet power supplies on the Right Arm. Modest current running should be possible, the problems are more severe when the rear of the hall gets a large dose during beam quality work.